OpenStudy (anonymous):

How do I write the matrix [1 0 0 | 0] [0 1 0 | 0] [0 0 0 | 0] as a basic solution? or parametric form? (This is a 3x3 matrix btw)

OpenStudy (anonymous):

Do you know what basic solution or parametric form would look like? Not for this particular matrix, but in general.

OpenStudy (anonymous):

yes :D, I'm just confused but the column of 0's

OpenStudy (anonymous):

Can you give me the general forms, because I don't completely remember them.

OpenStudy (anonymous):

I mean, I know how to simplify it, just not the names for it.

OpenStudy (anonymous):

\[ x=0\\ y=0\\ z=t \]Basically, \(x\) and \(y\) must be \(0\), while \(z\) can be whatever you want it to be.

OpenStudy (anonymous):

Another way to write this is: \[ \begin{bmatrix} 0\\0\\1 \end{bmatrix}t \]

OpenStudy (anonymous):

parametric is like: (An example) x1 = (9/5)s - (14/5)t x2 = (-4/5)s - (1/5)t x3 = s x4 = t basic solution is like: x1[s] = [9/5] [-4/5] [1] [0} x2[t] = [-14/5] [-1/5] [0] [1]

OpenStudy (anonymous):

yep just like that xD

OpenStudy (anonymous):

thank u :D! when you write the x=[0] y=[0] z=[t] solution, do you go according to row or column of the matrix ?

OpenStudy (anonymous):

For any row of \(0\)s, I would assign to \(t,u,v,\ldots\)

OpenStudy (anonymous):

So the first thing I did was say \(z=t\).

OpenStudy (anonymous):

Now, the \(x\) row basically turned into: \[ 1x+0y+0z = 0 \implies x=0 \]

OpenStudy (anonymous):

ahhhh i understand! so next row would be: 0x + 1y + 0z = 0, y = 0

OpenStudy (anonymous):

If the top row had been something like [1 0 2 | 4], then I would have gotten: \[ 1x+0y+2z = 4\implies x = 4-2z \implies x=4-2t \]

OpenStudy (anonymous):

ah i see. gotcha! thanks a bunch mate :D !

OpenStudy (anonymous):

Now, if the top row was something line [1 1 1 | 1]

OpenStudy (anonymous):

There would be a \(y\) term, which would seem problematic.

OpenStudy (anonymous):

But, I would have solved for \(y\) in the second column. So I could just substitute what I got for that in to the \(y\) term.

OpenStudy (anonymous):

But yeah, do a more advanced problem and it will make more sense.

OpenStudy (anonymous):

definitely! thank you so much :D

OpenStudy (anonymous):

The only real trick to this is to convert from: \[ \begin{bmatrix} 1&0&0\\ 0&1&0\\ 0&0&0 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix}= \begin{bmatrix} 0\\0\\0 \end{bmatrix} \]to \[ \begin{bmatrix} 1&0&0\\ 0&1&0\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix}= \begin{bmatrix} 0\\0\\t \end{bmatrix} \]Then it ends up being really simple.

OpenStudy (anonymous):

OHHHH i see it

OpenStudy (anonymous):

okay so for example [1 1 0 | 0] [0 0 1 | 0] [0 0 0 | 0] It would be [-1] t [0] [1] and if t = 1, then it would just be the same [-1] [0] [1]!

OpenStudy (anonymous):

Hold on...

OpenStudy (anonymous):

Our matrix is: \[ \begin{bmatrix} 1&1&0\\ 0&0&1\\ 0&0&0 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix}= \begin{bmatrix} 0\\0\\0 \end{bmatrix} \]We can swap rows:\[ \begin{bmatrix} 1&1&0\\ 0&0&0\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix}= \begin{bmatrix} 0\\0\\0 \end{bmatrix} \]Then we plug in the free variable:\[ \begin{bmatrix} 1&1&0\\ 0&1&0\\ 0&0&1 \end{bmatrix} \begin{bmatrix} x\\y\\z \end{bmatrix}= \begin{bmatrix} 0\\t\\0 \end{bmatrix} \]Which gives us the equations:\[ x+y=0\implies x+t=0\implies x=-t\\ y=t\\ z=0 \]

OpenStudy (anonymous):

\[ \begin{bmatrix} x\\y\\z \end{bmatrix} = \begin{bmatrix} -1\\1\\0 \end{bmatrix}t \]

OpenStudy (anonymous):

oh crap, ah i see. you did the exact same thing you wrote above

OpenStudy (anonymous):

I found it a bit confusing at first as well. They key thing to remember is that these are just systems of equations written in a weird way.

OpenStudy (anonymous):

Want to see something interesting?

OpenStudy (anonymous):

ahhh omg i see. and sure!

OpenStudy (anonymous):

i'm doing eigenvectors and eigenvalues and it's just like wow i forgot how to do basic matrices lol

OpenStudy (anonymous):

\[ \begin{bmatrix} 0&0\\ -m&1 \end{bmatrix} \begin{bmatrix} x\\y \end{bmatrix} = \begin{bmatrix} 0\\b \end{bmatrix} \]Suppose we just let the free variable for \(x\) be \(x\).\[ \begin{bmatrix} 1&0\\ -m&1 \end{bmatrix} \begin{bmatrix} x\\y \end{bmatrix} = \begin{bmatrix} x\\b \end{bmatrix} \]This gives us equations: \[ x=x\\ -mx+y=b\implies y=mx+b \]

OpenStudy (anonymous):

Our line is just a system of equations, where \(x\) is a free variable.

OpenStudy (anonymous):

i never thought of it like this with matrices omfg it's true

OpenStudy (anonymous):

I think you can do the rest. The only tricks here are to swap your rows so that their first non-zero term is in the diagonal, then put in free variables.

OpenStudy (anonymous):

thank you so much ^^

OpenStudy (anonymous):

Or just recognize that if a column doesn't have a pivot, it's a free variable. You're welcome.